# Descent-Guided Policy Gradient for Scalable Cooperative Multi-Agent Learning

### v1

@@ E.1 Lemma: Guidance Consistency
Under Assumptions 3.1 and 3.2, if $\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta}^{*})=\mathbf{0}$, then $\nabla_{\boldsymbol{\theta}}\mathcal{G}(\boldsymbol{\theta}^{*})=\mathbf{0}$.
@@
We proceed by contradiction. Let $\boldsymbol{x}^{*}\triangleq\boldsymbol{x}(\boldsymbol{\theta}^{*})$ denote the system state induced by the stationary point $\boldsymbol{\theta}^{*}$, and suppose $\nabla_{\boldsymbol{\theta}}\mathcal{G}(\boldsymbol{\theta}^{*})\neq\mathbf{0}$.
Step 1. If $\boldsymbol{x}^{*}=\tilde{\boldsymbol{x}}$, then $\mathcal{G}(\boldsymbol{\theta}^{*})=\mathbb{E}[\sum_{t}\gamma^{t}\tfrac{1}{2}\|\tilde{\boldsymbol{x}}_{t}-\tilde{\boldsymbol{x}}_{t}\|^{2}]=0$, which is the global minimum of $\mathcal{G}$, and hence $\nabla_{\boldsymbol{\theta}}\mathcal{G}(\boldsymbol{\theta}^{*})=\mathbf{0}$—contradicting our assumption. Therefore $\boldsymbol{x}^{*}\neq\tilde{\boldsymbol{x}}$.
Step 2. By Assumption 3.2, for any state not equal to the reference:
$\boldsymbol{x}^{*}\neq\tilde{\boldsymbol{x}}\implies\langle\nabla_{\boldsymbol{x}}J(\boldsymbol{x}^{*}),\,\tilde{\boldsymbol{x}}-\boldsymbol{x}^{*}\rangle>0.$
The right-hand side is strictly positive, and $\tilde{\boldsymbol{x}}-\boldsymbol{x}^{*}\neq\mathbf{0}$, so $\nabla_{\boldsymbol{x}}J(\boldsymbol{x}^{*})\neq\mathbf{0}$.
Step 3. Under sufficient policy expressiveness (see Remark below), a non-zero state-space gradient lifts to a non-zero parameter-space gradient: $\nabla_{\boldsymbol{x}}J(\boldsymbol{x}^{*})\neq\mathbf{0}$ implies $\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta}^{*})\neq\mathbf{0}$. This contradicts the premise $\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta}^{*})=\mathbf{0}$.
Therefore $\nabla_{\boldsymbol{\theta}}\mathcal{G}(\boldsymbol{\theta}^{*})=\mathbf{0}$. ∎
@@
The proof requires the Jacobian $\nabla_{\boldsymbol{\theta}}\boldsymbol{x}$ to have full row rank, ensuring that every non-zero state-space gradient $\nabla_{\boldsymbol{x}}J$ lifts to a non-zero parameter-space gradient $\nabla_{\boldsymbol{\theta}}J$. This is a standard condition for overparameterized neural network policies ($p\gg d$) and holds generically in the absence of degenerate symmetries.
@@
At a stationary point $\boldsymbol{\theta}^{*}$ of $J$, we have $\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta}^{*})=\mathbf{0}$. By Lemma E.1, $\nabla_{\boldsymbol{\theta}}\mathcal{G}(\boldsymbol{\theta}^{*})=\mathbf{0}$. Therefore:
$\nabla_{\boldsymbol{\theta}}J_{\alpha}(\boldsymbol{\theta}^{*})=(1-\alpha)\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta}^{*})-\alpha\nabla_{\boldsymbol{\theta}}\mathcal{G}(\boldsymbol{\theta}^{*})=\mathbf{0},$
confirming that every stationary point of the original objective $J$ is also a stationary point of the regularized objective $J_{\alpha}$. ∎
@@
In this appendix, we prove that the DG-PG estimator achieves agent-independent variance scaling. As established in Section 4.2, the standard policy gradient estimator has variance $\sigma_{J}^{2}=\Theta(N)$ due to cross-agent exploration noise (Kuba et al., 2021), while the guidance gradient estimator has variance $\sigma_{\mathcal{G}}^{2}=\mathcal{O}(1)$ since its coefficient depends only on deterministic quantities given a realized trajectory.
Theorem 4.2 (Restated). Under Assumptions 3.13.2, the DG-PG estimator achieves $\sigma_{\text{DG}}^{2}=\mathcal{O}(1)$ at the optimal mixing coefficient, independent of the number of agents $N$.
@@
$\sigma_{\text{DG}}^{2}(\alpha)=(1-\alpha)^{2}\sigma_{J}^{2}+\alpha^{2}\sigma_{\mathcal{G}}^{2}-2\alpha(1-\alpha)\rho\sigma_{J}\sigma_{\mathcal{G}},$
where $\rho=\text{Corr}(\hat{g}_{J}^{i},\hat{g}_{\mathcal{G}}^{i})$. By Assumption 3.2, the guidance direction is aligned with improving $J$, ensuring $0<|\rho|<1$.
Since $\sigma_{\text{DG}}^{2}(\alpha)$ is a convex quadratic in $\alpha$, setting $\frac{d}{d\alpha}\sigma_{\text{DG}}^{2}=0$ yields the optimal mixing coefficient:
$\alpha^{*}=\frac{\sigma_{J}^{2}+\rho\,\sigma_{J}\sigma_{\mathcal{G}}}{\sigma_{J}^{2}+\sigma_{\mathcal{G}}^{2}+2\rho\,\sigma_{J}\sigma_{\mathcal{G}}}.$
@@
$\sigma_{\text{DG},\min}^{2}=\frac{\sigma_{J}^{2}\sigma_{\mathcal{G}}^{2}(1-\rho^{2})}{\sigma_{J}^{2}+\sigma_{\mathcal{G}}^{2}+2\rho\sigma_{J}\sigma_{\mathcal{G}}}.$
As $N\to\infty$, we have $\sigma_{J}^{2}=\Theta(N)$ and $\sigma_{\mathcal{G}}^{2}=\mathcal{O}(1)$. Dividing numerator and denominator by $\sigma_{J}^{2}$:
$\sigma_{\text{DG},\min}^{2}=\frac{\sigma_{\mathcal{G}}^{2}(1-\rho^{2})}{1+\mathcal{O}(N^{-1})}\xrightarrow{N\to\infty}\sigma_{\mathcal{G}}^{2}(1-\rho^{2})=\mathcal{O}(1).$
@@
The optimal mixing coefficient $\alpha^{*}\to 1$ as $N\to\infty$, indicating that large-scale systems should rely primarily on the guidance signal to suppress cross-agent variance. In practice, we adopt a dynamic $\alpha$ schedule—starting with a large $\alpha$ for fast early convergence and decaying to a smaller value to allow the RL signal to fine-tune beyond the reference (see Section 6).
@@
Theorem 4.3 (Restated). Under the smoothness and PL conditions (Assumption G.1 below) and with $\sigma_{\text{DG}}^{2}=\mathcal{O}(1)$ from Theorem 4.2, DG-PG achieves an $\epsilon$-optimal policy in $T=\mathcal{O}(L/(\mu\epsilon))$ iterations, independent of $N$.
@@
The regularized objective $J_{\alpha}(\boldsymbol{\theta})$ is $L$-smooth and satisfies the Polyak-Łojasiewicz condition with parameter $\mu>0$:
$\frac{1}{2}\|\nabla J_{\alpha}(\boldsymbol{\theta})\|^{2}\geq\mu(J_{\alpha}^{*}-J_{\alpha}(\boldsymbol{\theta})).$
These conditions are standard in the stochastic optimization literature (Karimi et al., 2016); they are needed only to convert the $\mathcal{O}(1)$ variance bound (Theorem 4.2)—which holds unconditionally—into an explicit iteration complexity. Under weaker regularity (e.g., smoothness without PL), the rate changes but the $N$-independence is preserved.
@@
The following analysis holds for a fixed mixing parameter $\alpha\in(0,1)$, so that $J_{\alpha}$ is a fixed objective with well-defined smoothness and PL constants. When $\alpha$ is varied dynamically in practice, the convergence guarantee applies to each phase as a snapshot analysis.
Let $\sigma_{\text{DG}}^{2}\triangleq\mathrm{Var}(\hat{g}_{\text{DG}})=\mathcal{O}(1)$ by Theorem 4.2. Consider the stochastic gradient ascent update
$\boldsymbol{\theta}_{k+1}=\boldsymbol{\theta}_{k}+\eta\hat{g}_{\text{DG}}(\boldsymbol{\theta}_{k})$
with step size $\eta\leq\frac{1}{L}$.
By $L$-smoothness of $J_{\alpha}$:
$J_{\alpha}(\boldsymbol{\theta}_{k+1})\geq J_{\alpha}(\boldsymbol{\theta}_{k})+\langle\nabla J_{\alpha}(\boldsymbol{\theta}_{k}),\,\boldsymbol{\theta}_{k+1}-\boldsymbol{\theta}_{k}\rangle-\frac{L}{2}\|\boldsymbol{\theta}_{k+1}-\boldsymbol{\theta}_{k}\|^{2}.$
@@
$\mathbb{E}_{k}[J_{\alpha}(\boldsymbol{\theta}_{k+1})]\geq J_{\alpha}(\boldsymbol{\theta}_{k})+\eta\langle\nabla J_{\alpha}(\boldsymbol{\theta}_{k}),\,\mathbb{E}_{k}[\hat{g}_{\text{DG}}]\rangle-\frac{L\eta^{2}}{2}\mathbb{E}_{k}[\|\hat{g}_{\text{DG}}\|^{2}].$
Since $\hat{g}_{\text{DG}}$ is an unbiased estimator of $\nabla J_{\alpha}$, we have $\mathbb{E}_{k}[\hat{g}_{\text{DG}}]=\nabla J_{\alpha}(\boldsymbol{\theta}_{k})$ and $\mathbb{E}_{k}[\|\hat{g}_{\text{DG}}\|^{2}]=\|\nabla J_{\alpha}(\boldsymbol{\theta}_{k})\|^{2}+\sigma_{\text{DG}}^{2}$. Substituting:
$\mathbb{E}_{k}[J_{\alpha}(\boldsymbol{\theta}_{k+1})]$ $\geq J_{\alpha}(\boldsymbol{\theta}_{k})+\eta\|\nabla J_{\alpha}(\boldsymbol{\theta}_{k})\|^{2}-\frac{L\eta^{2}}{2}\left(\|\nabla J_{\alpha}(\boldsymbol{\theta}_{k})\|^{2}+\sigma_{\text{DG}}^{2}\right)$
$=J_{\alpha}(\boldsymbol{\theta}_{k})+\left(\eta-\frac{L\eta^{2}}{2}\right)\|\nabla J_{\alpha}(\boldsymbol{\theta}_{k})\|^{2}-\frac{L\eta^{2}}{2}\sigma_{\text{DG}}^{2}.$
Since $\eta\leq\frac{1}{L}$, we have $\eta-\frac{L\eta^{2}}{2}\geq\frac{\eta}{2}$, giving:
@@
$\mathbb{E}_{k}[J_{\alpha}(\boldsymbol{\theta}_{k+1})]\geq J_{\alpha}(\boldsymbol{\theta}_{k})+\frac{\eta}{2}\|\nabla J_{\alpha}(\boldsymbol{\theta}_{k})\|^{2}-\frac{L\eta^{2}}{2}\sigma_{\text{DG}}^{2}.$
Subtracting both sides of the above inequality from $J_{\alpha}^{*}$ and taking full expectation flips the direction:
@@
$J_{\alpha}^{*}-\mathbb{E}[J_{\alpha}(\boldsymbol{\theta}_{k+1})]$ $\leq J_{\alpha}^{*}-J_{\alpha}(\boldsymbol{\theta}_{k})-\frac{\eta}{2}\|\nabla J_{\alpha}(\boldsymbol{\theta}_{k})\|^{2}+\frac{L\eta^{2}}{2}\sigma_{\text{DG}}^{2}.$
Let $\Delta_{k}\triangleq\mathbb{E}[J_{\alpha}^{*}-J_{\alpha}(\boldsymbol{\theta}_{k})]$ denote the expected optimality gap. In terms of $\Delta_{k}$:
@@
$\Delta_{k+1}\leq\Delta_{k}-\frac{\eta}{2}\|\nabla J_{\alpha}(\boldsymbol{\theta}_{k})\|^{2}+\frac{L\eta^{2}}{2}\sigma_{\text{DG}}^{2}.$
By the PL condition (Assumption G.1), $\|\nabla J_{\alpha}(\boldsymbol{\theta}_{k})\|^{2}\geq 2\mu(J_{\alpha}^{*}-J_{\alpha}(\boldsymbol{\theta}_{k}))=2\mu\Delta_{k}$. Substituting:
$\Delta_{k+1}$ $\leq\Delta_{k}-\frac{\eta}{2}\cdot 2\mu\Delta_{k}+\frac{L\eta^{2}}{2}\sigma_{\text{DG}}^{2}$

### v2

@@ E.1 Lemma: Guidance Consistency
Under Assumptions 3.1 and 3.2, if $\nabla_{\theta^{i}}J(\boldsymbol{\theta}^{*})=\mathbf{0}$ for all $i\in\mathcal{N}$, then $\nabla_{\theta^{i}}\mathcal{G}(\boldsymbol{\theta}^{*})=\mathbf{0}$ for all $i\in\mathcal{N}$.
@@
We proceed by contradiction. Suppose $\nabla_{\theta^{i}}J(\boldsymbol{\theta}^{*})=\mathbf{0}$ for all $i\in\mathcal{N}$ (the Nash condition), but $\nabla_{\theta^{i}}\mathcal{G}(\boldsymbol{\theta}^{*})\neq\mathbf{0}$ for some agent $i$. Recall that $\mathcal{G}(\boldsymbol{\theta})=\mathbb{E}_{s\sim\nu^{\boldsymbol{\pi}}}[d(\boldsymbol{x}^{\boldsymbol{\pi}}|_{s},\tilde{\boldsymbol{x}}|_{s})]$ is the expected deviation under the state visitation distribution $\nu^{\boldsymbol{\pi}}$, where $\boldsymbol{x}^{\boldsymbol{\pi}}|_{s}$ is the system state jointly determined by agents’ actions under policy $\boldsymbol{\pi}$ at MDP state $s$, and $\tilde{\boldsymbol{x}}|_{s}$ is the corresponding reference. We write $\boldsymbol{x}^{*}\triangleq\boldsymbol{x}^{\boldsymbol{\pi}^{*}}|_{s}$ and $\tilde{\boldsymbol{x}}\triangleq\tilde{\boldsymbol{x}}|_{s}$ for a given MDP state $s$, where $\boldsymbol{\pi}^{*}\triangleq\boldsymbol{\pi}_{\boldsymbol{\theta}^{*}}$.
Step 1. If $\boldsymbol{x}^{*}=\tilde{\boldsymbol{x}}$ for all states $s$ visited under $\nu^{\boldsymbol{\pi}^{*}}$, then $d(\boldsymbol{x}^{*},\tilde{\boldsymbol{x}})=0$ everywhere and $\mathcal{G}(\boldsymbol{\theta}^{*})=0$. This is the global minimum of $\mathcal{G}\geq 0$, so the first-order condition gives $\nabla_{\theta^{i}}\mathcal{G}(\boldsymbol{\theta}^{*})=\mathbf{0}$ for all $i$, contradicting our assumption. Therefore there exists a visited state $s$ at which $\boldsymbol{x}^{*}\neq\tilde{\boldsymbol{x}}$.
Step 2. By Assumption 3.2, at any state $s$ where $\boldsymbol{x}^{*}\neq\tilde{\boldsymbol{x}}$:
$\langle\nabla_{\boldsymbol{x}}J(\boldsymbol{x}^{*}),\,\tilde{\boldsymbol{x}}-\boldsymbol{x}^{*}\rangle>0.$
Since $\tilde{\boldsymbol{x}}-\boldsymbol{x}^{*}\neq\mathbf{0}$, the state-space gradient $\nabla_{\boldsymbol{x}}J(\boldsymbol{x}^{*})$ must be non-zero.
Step 3. Under sufficient policy expressiveness (see Remark E.2), a non-zero state-space gradient $\nabla_{\boldsymbol{x}}J(\boldsymbol{x}^{*})\neq\mathbf{0}$ implies $\nabla_{\theta^{i}}J(\boldsymbol{\theta}^{*})\neq\mathbf{0}$ for some agent $i$. This contradicts the Nash condition.
Therefore $\nabla_{\theta^{i}}\mathcal{G}(\boldsymbol{\theta}^{*})=\mathbf{0}$ for all $i\in\mathcal{N}$. ∎
@@
The proof requires the Jacobian $\nabla_{\theta^{i}}\boldsymbol{x}$ to have full row rank for at least one agent $i$, ensuring that a non-zero state-space gradient $\nabla_{\boldsymbol{x}}J$ lifts to a non-zero parameter-space gradient $\nabla_{\theta^{i}}J$. This is a standard condition for overparameterized neural network policies ($p\gg m$) and holds generically in the absence of degenerate symmetries.
@@
At a Nash equilibrium $\boldsymbol{\theta}^{*}$, we have $\nabla_{\theta^{i}}J(\boldsymbol{\theta}^{*})=\mathbf{0}$ for all $i\in\mathcal{N}$. By Lemma E.1, $\nabla_{\theta^{i}}\mathcal{G}(\boldsymbol{\theta}^{*})=\mathbf{0}$ for all $i\in\mathcal{N}$. Therefore:
$\nabla_{\theta^{i}}J_{\alpha}(\boldsymbol{\theta}^{*})=(1-\alpha)\nabla_{\theta^{i}}J(\boldsymbol{\theta}^{*})-\alpha\nabla_{\theta^{i}}\mathcal{G}(\boldsymbol{\theta}^{*})=\mathbf{0},$
confirming that the augmentation does not alter the Nash equilibria of the original game. ∎
@@
In this appendix, we prove that the DG-PG estimator achieves agent-independent variance scaling. As established in Section 4.2, the standard policy gradient estimator has variance $\sigma_{J}^{2}=\Theta(N)$ due to cross-agent exploration noise (Kuba et al., 2021), while the gradient estimator for the guidance term has variance $\sigma_{\mathcal{G}}^{2}=\mathcal{O}(1)$ since its coefficient depends only on deterministic quantities given a realized trajectory.
Theorem 4.2 (Agent-Independent Variance Bound). Under Assumptions 3.13.2, let $\sigma_{J}^{2}\triangleq\mathrm{Var}(\hat{g}_{J}^{i})$, $\sigma_{\mathcal{G}}^{2}\triangleq\mathrm{Var}(\hat{g}_{\mathcal{G}}^{i})$, and $\rho\triangleq\mathrm{Corr}(\hat{g}_{J}^{i},\hat{g}_{\mathcal{G}}^{i})\in(-1,0)$. The optimally weighted combined estimator achieves:
$\min_{\alpha\in(0,1)}\mathrm{Var}(\hat{g}_{\mathrm{DG}}^{i})=\frac{\sigma_{J}^{2}\,\sigma_{\mathcal{G}}^{2}\,(1-\rho^{2})}{\sigma_{J}^{2}+\sigma_{\mathcal{G}}^{2}+2\rho\,\sigma_{J}\sigma_{\mathcal{G}}}.$
Since $\sigma_{J}^{2}=\Theta(N)$ and $\sigma_{\mathcal{G}}^{2}=\mathcal{O}(1)$, the minimum variance satisfies $\min_{\alpha}\mathrm{Var}(\hat{g}_{\mathrm{DG}}^{i})=\mathcal{O}(1)$ for all $N$, independent of the number of agents.
@@
$\sigma_{\text{DG}}^{2}(\alpha)=(1-\alpha)^{2}\sigma_{J}^{2}+\alpha^{2}\sigma_{\mathcal{G}}^{2}-2\alpha(1-\alpha)\rho\sigma_{J}\sigma_{\mathcal{G}},$ (20)
where $\rho=\text{Corr}(\hat{g}_{J}^{i},\hat{g}_{\mathcal{G}}^{i})\in(-1,0)$ by Assumption 3.2.
The combined variance $\sigma_{\text{DG}}^{2}(\alpha)$ is a convex quadratic in $\alpha$. To minimize it, we set $\frac{d}{d\alpha}\sigma_{\text{DG}}^{2}=0$:
$-2(1-\alpha)\sigma_{J}^{2}+2\alpha\sigma_{\mathcal{G}}^{2}-2(1-2\alpha)\rho\sigma_{J}\sigma_{\mathcal{G}}=0\quad\Longrightarrow\quad\alpha^{*}=\frac{\sigma_{J}^{2}+\rho\,\sigma_{J}\sigma_{\mathcal{G}}}{\sigma_{J}^{2}+\sigma_{\mathcal{G}}^{2}+2\rho\,\sigma_{J}\sigma_{\mathcal{G}}}.$ (21)
@@
$\sigma_{\text{DG},\min}^{2}=\frac{\sigma_{J}^{2}\sigma_{\mathcal{G}}^{2}(1-\rho^{2})}{\sigma_{J}^{2}+\sigma_{\mathcal{G}}^{2}+2\rho\sigma_{J}\sigma_{\mathcal{G}}}.$ (22)
As established in Section 4.2, $\sigma_{J}^{2}=\Theta(N)$ because the return aggregates all $N$ agents’ stochastic actions, while $\sigma_{\mathcal{G}}^{2}=\mathcal{O}(1)$ because the guidance coefficient is computed from deterministic quantities. Dividing numerator and denominator by $\sigma_{J}^{2}$:
$\sigma_{\text{DG},\min}^{2}=\frac{\sigma_{\mathcal{G}}^{2}(1-\rho^{2})}{1+\frac{\sigma_{\mathcal{G}}^{2}}{\sigma_{J}^{2}}+2\rho\frac{\sigma_{\mathcal{G}}}{\sigma_{J}}}=\frac{\sigma_{\mathcal{G}}^{2}(1-\rho^{2})}{1+\mathcal{O}(N^{-1})}.$
The numerator $\sigma_{\mathcal{G}}^{2}(1-\rho^{2})=\mathcal{O}(1)$ since $\sigma_{\mathcal{G}}^{2}=\mathcal{O}(1)$ and $\rho\in(-1,0)$ is a constant. The denominator $1+\mathcal{O}(N^{-1})$ is bounded away from zero for all $N\geq 1$. Therefore:
$\sigma_{\text{DG},\min}^{2}=\mathcal{O}(1),\quad\text{for all }N\geq 1.$ (23)
@@
The optimal mixing coefficient $\alpha^{*}\to 1$ as $N\to\infty$, indicating that large-scale systems should rely primarily on the guidance signal to suppress cross-agent variance. In practice, we adopt a dynamic $\alpha$ schedule, starting with a large $\alpha$ for fast early convergence and decaying to a smaller value to allow the RL signal to fine-tune beyond the reference (see Section 6).
@@
Theorem 4.3 (Sample Complexity). Under the smoothness and PL conditions (Assumption G.1 below) and with $\sigma_{\text{DG}}^{2}=\mathcal{O}(1)$ from Theorem 4.2, DG-PG achieves an $\epsilon$-optimal policy in $T=\mathcal{O}\bigl(\frac{L}{\mu^{2}\epsilon}\log\frac{1}{\epsilon}\bigr)$ iterations, independent of $N$.
@@
The cooperative objective $J(\boldsymbol{\theta})$ is $L$-smooth and satisfies the Polyak-Łojasiewicz (PL) condition with PL constant $\mu\in(0,L]$:
$\frac{1}{2}\|\nabla J(\boldsymbol{\theta})\|^{2}\geq\mu(J^{*}-J(\boldsymbol{\theta})),$
where $J^{*}\triangleq\sup_{\boldsymbol{\theta}}J(\boldsymbol{\theta})$ denotes the optimal value of the cooperative objective.
Together, these conditions guarantee that stochastic gradient ascent achieves linear convergence of the optimality gap. These are standard conditions in stochastic optimization (Karimi et al., 2016), needed only to convert the $\mathcal{O}(1)$ variance bound (Theorem 4.2) into an explicit iteration complexity. They are commonly satisfied for smooth policy parameterizations such as softmax policies (Mei et al., 2020; Agarwal et al., 2021).
@@
Our proof first characterizes, through $L$-smoothness and the update rule, how the variance and bias of $\hat{g}_{\text{DG}}$ affect the per-iteration change in $J$. We then apply the PL condition to connect the gradient norm $\|\nabla J\|^{2}$ to the optimality gap $J^{*}-J(\boldsymbol{\theta}_{k})$, thereby obtaining the $N$-independent iteration complexity. The following analysis holds for a fixed mixing parameter $\alpha\in(0,1)$. When $\alpha$ is varied dynamically in practice, the convergence guarantee applies to each phase as a snapshot analysis.
Step 1: Smoothness bound. Let $\bar{g}_{\text{DG}}(\boldsymbol{\theta}_{k})\triangleq\mathbb{E}_{k}[\hat{g}_{\text{DG}}(\boldsymbol{\theta}_{k})]$ denote the expected augmented gradient. Consider the stochastic gradient ascent update
$\boldsymbol{\theta}_{k+1}=\boldsymbol{\theta}_{k}+\eta\hat{g}_{\text{DG}}(\boldsymbol{\theta}_{k})$ (24)
By $L$-smoothness of $J$:
$J(\boldsymbol{\theta}_{k+1})\geq J(\boldsymbol{\theta}_{k})+\langle\nabla J(\boldsymbol{\theta}_{k}),\,\boldsymbol{\theta}_{k+1}-\boldsymbol{\theta}_{k}\rangle-\frac{L}{2}\|\boldsymbol{\theta}_{k+1}-\boldsymbol{\theta}_{k}\|^{2}.$
@@
$\mathbb{E}_{k}[J(\boldsymbol{\theta}_{k+1})]\geq J(\boldsymbol{\theta}_{k})+\eta\langle\nabla J(\boldsymbol{\theta}_{k}),\,\bar{g}_{\text{DG}}(\boldsymbol{\theta}_{k})\rangle-\frac{L\eta^{2}}{2}\mathbb{E}_{k}[\|\hat{g}_{\text{DG}}(\boldsymbol{\theta}_{k})\|^{2}].$ (25)
Step 2: Per-iteration descent bound. The descent inequality Equation 25 shows that the per-iteration ascent on $J$ is governed by two quantities: the inner product term $\langle\nabla J,\bar{g}_{\text{DG}}\rangle$, which captures the alignment between the true gradient and the augmented gradient, and the second moment term $\mathbb{E}_{k}[\|\hat{g}_{\text{DG}}\|^{2}]$, which captures the estimator’s variance and bias. We bound both independently of $N$.
For the inner product term $\langle\nabla J(\boldsymbol{\theta}_{k}),\,\bar{g}_{\text{DG}}(\boldsymbol{\theta}_{k})\rangle$, by definition of the expected augmented gradient:
$\bar{g}_{\text{DG}}(\boldsymbol{\theta}_{k})=(1-\alpha)\nabla J(\boldsymbol{\theta}_{k})-\alpha\nabla\mathcal{G}(\boldsymbol{\theta}_{k}).$
Therefore:
@@
$\langle\nabla J(\boldsymbol{\theta}_{k}),\,\bar{g}_{\text{DG}}(\boldsymbol{\theta}_{k})\rangle=(1-\alpha)\|\nabla J(\boldsymbol{\theta}_{k})\|^{2}-\alpha\langle\nabla J(\boldsymbol{\theta}_{k}),\,\nabla\mathcal{G}(\boldsymbol{\theta}_{k})\rangle.$
Since improving $J$ reduces the deviation $\mathcal{G}$ (Assumption 3.2), the true gradients satisfy $\langle\nabla J(\boldsymbol{\theta}_{k}),\nabla\mathcal{G}(\boldsymbol{\theta}_{k})\rangle\leq 0$, so $\alpha\langle\nabla J,\nabla\mathcal{G}\rangle\geq 0$. Dropping this non-negative term yields the lower bound for the inner product term in Equation 25:
$\langle\nabla J(\boldsymbol{\theta}_{k}),\,\bar{g}_{\text{DG}}(\boldsymbol{\theta}_{k})\rangle\geq(1-\alpha)\|\nabla J(\boldsymbol{\theta}_{k})\|^{2}.$ (26)
For the second moment $\mathbb{E}_{k}[\|\hat{g}_{\text{DG}}(\boldsymbol{\theta}_{k})\|^{2}]$, by the bias-variance decomposition:
$\mathbb{E}_{k}[\|\hat{g}_{\text{DG}}(\boldsymbol{\theta}_{k})\|^{2}]=\|\bar{g}_{\text{DG}}(\boldsymbol{\theta}_{k})\|^{2}+\sigma_{\text{DG}}^{2}.$ (27)
The variance $\sigma_{\text{DG}}^{2}=\mathcal{O}(1)$ by Theorem 4.2. Since $\bar{g}_{\text{DG}}=(1-\alpha)\nabla J-\alpha\nabla\mathcal{G}$, applying the triangle inequality:
$\|\bar{g}_{\text{DG}}(\boldsymbol{\theta}_{k})\|^{2}\leq 2(1-\alpha)^{2}\|\nabla J(\boldsymbol{\theta}_{k})\|^{2}+2\alpha^{2}\|\nabla\mathcal{G}(\boldsymbol{\theta}_{k})\|^{2}.$ (28)
For the $\|\nabla\mathcal{G}(\boldsymbol{\theta}_{k})\|$ term, by Equation 9 the gradient of the guidance term has the form $\nabla_{\theta^{i}}\mathcal{G}=\mathbb{E}[\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\boldsymbol{z}_{t}^{i}\rangle\nabla_{\theta^{i}}\log\pi^{i}]$. Each factor is bounded:
  • $\|\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t}\|$ is bounded by compactness of the system state space.
  • $\|\boldsymbol{z}_{t}^{i}\|$ is bounded because each agent has bounded marginal effect on the system state.
  • $\|\nabla_{\theta^{i}}\log\pi^{i}\|$ is bounded by standard score function regularity.
Applying Cauchy–Schwarz to the inner product $\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\boldsymbol{z}_{t}^{i}\rangle$ and combining with the score function bound, there exists a constant $B>0$, independent of $N$, such that:
$\|\nabla\mathcal{G}(\boldsymbol{\theta}_{k})\|\leq B.$ (29)
Combining Equations 29 and 28 with Equation 27, we obtain:
$\mathbb{E}_{k}[\|\hat{g}_{\text{DG}}(\boldsymbol{\theta}_{k})\|^{2}]\leq 2(1-\alpha)^{2}\|\nabla J(\boldsymbol{\theta}_{k})\|^{2}+2\alpha^{2}B^{2}+\sigma_{\text{DG}}^{2}.$ (30)
which is the upper bound for the second moment term in Equation 25.
Substituting the inner product bound (Equation 26) and the second moment bound (Equation 30) into the descent inequality (Equation 25), we obtain the per-iteration descent bound:
@@
$\mathbb{E}_{k}[J(\boldsymbol{\theta}_{k+1})]$ $\geq J(\boldsymbol{\theta}_{k})+\eta(1-\alpha)\|\nabla J(\boldsymbol{\theta}_{k})\|^{2}-\frac{L\eta^{2}}{2}\left(2(1-\alpha)^{2}\|\nabla J(\boldsymbol{\theta}_{k})\|^{2}+2\alpha^{2}B^{2}+\sigma_{\text{DG}}^{2}\right)$
@@
$=J(\boldsymbol{\theta}_{k})+\left[\eta(1-\alpha)-L\eta^{2}(1-\alpha)^{2}\right]\|\nabla J(\boldsymbol{\theta}_{k})\|^{2}-\frac{L\eta^{2}}{2}(2\alpha^{2}B^{2}+\sigma_{\text{DG}}^{2}).$
Under the update rule (Equation 24), for the gradient signal to contribute positively to the expected change in $J$, the coefficient of $\|\nabla J(\boldsymbol{\theta}_{k})\|^{2}$ must be positive. Choosing $\eta\leq\frac{1}{2L(1-\alpha)}$ ensures $\eta(1-\alpha)-L\eta^{2}(1-\alpha)^{2}\geq\frac{\eta(1-\alpha)}{2}$:
$\mathbb{E}_{k}[J(\boldsymbol{\theta}_{k+1})]\geq J(\boldsymbol{\theta}_{k})+\frac{\eta(1-\alpha)}{2}\|\nabla J(\boldsymbol{\theta}_{k})\|^{2}-\frac{L\eta^{2}}{2}(2\alpha^{2}B^{2}+\sigma_{\text{DG}}^{2}).$ (31)
Step 3: PL condition and recursion. We now use the PL condition to convert the gradient norm bound into optimality gap contraction.
Subtracting both sides of Equation 31 from $J^{*}$ (Assumption G.1):
$J^{*}-\mathbb{E}_{k}[J(\boldsymbol{\theta}_{k+1})]\leq J^{*}-J(\boldsymbol{\theta}_{k})-\frac{\eta(1-\alpha)}{2}\|\nabla J(\boldsymbol{\theta}_{k})\|^{2}+\frac{L\eta^{2}}{2}(2\alpha^{2}B^{2}+\sigma_{\text{DG}}^{2}).$
Taking full expectation and defining $\Delta_{k}\triangleq\mathbb{E}[J^{*}-J(\boldsymbol{\theta}_{k})]$: